New York Journal of Mathematics

Volume 1

Editor in Chief Mark Steinberger

Editors:

Marc Culler Andrew Granville Jonathan Lubin Paul Muhly Bodo Pareigis Andrew Ranicki Douglas C. Ravenel Richard Rochberg Joseph Rosenblatt David J. Saltman Birgit Speh Ronald J. Stern

The New York Journal of Mathematics is offered electronically, via the following modes of access.

• WWW to http://nyjm.albany.edu:8000/nyjm.html

• gopher to nyjm.albany.edu 1070

• anonymous ftp to nyjm.albany.edu, in /pub/nyjm Papers should be submitted by electronic mail to the editor closest to your field. Information regarding submissions is contained in the internet archive accessed by the above means.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994 by the State University of New York. ISSN 1076-9803